# The taggedPBC: Annotating a massive parallel corpus for crosslinguistic investigations

### v2

@@ Abstract
Existing datasets available for crosslinguistic investigations have tended to focus on large amounts of data for a small group of languages or a small amount of data for a large number of languages. This means that claims based on these datasets are limited in what they reveal about universal properties of the human language faculty. While this has begun to change through the efforts of projects seeking to develop tagged corpora for a large number of languages, such efforts are still constrained by limits on resources. The current paper reports on a large tagged parallel dataset which has been developed to partially address this issue. The taggedPBC contains POS-tagged parallel text data from more than 1,940 languages, representing 155 language families and 78 isolates, dwarfing previously available resources. The accuracy of particular tags in this dataset is shown to correlate well with both existing SOTA taggers for high-resource languages (SpaCy, Trankit) as well as hand-tagged corpora (Universal Dependencies Treebanks). Additionally, a novel measure derived from this dataset, the N1 ratio, correlates with expert determinations of intransitive word order in three typological databases (WALS, Grambank, Autotyp) such that a Gaussian Naive Bayes classifier trained on this feature can accurately identify basic intransitive word order for languages not in those databases. While much work is still needed to expand and develop this dataset, the taggedPBC is an important step to enable corpus-based crosslinguistic investigations, and is made available for research and collaboration via GitHub.
@@ 1 Introduction
There are over 7,000 languages in the world, and such languages are rather diverse in terms of their typological features. At the same time, despite the overall diversity there are many similarities between languages, suggesting the possibility of universal properties that guide language evolution. These diversities and similarities have been identified and categorized in various ways by a variety of databases, most notably WALS (Dryer & Haspelmath 2013), [^1] https://wals.info/ Grambank (Skirgård et al., 2023), [^2] https://grambank.clld.org/ Autotyp (Bickel et al., 2023), [^3] https://github.com/autotyp/autotyp-data Glottolog (Hammarström et al. 2024), [^4] https://www.glottolog.org and Lexibank (List et al. 2022). [^5] https://lexibank.clld.org/ There are also collections of corpora that are expansive in their coverage of individual or groups of languages yet focused on particular genres or domains (CHILDES, [^6] https://childes.talkbank.org/ Brown, etc; and see the extensive corpora available via the CLARIN [^7] https://www.clarin.eu/resource-families project). The former tend to provide information about languages, while the latter tend to provide actual language data in various forms. The issue with the former is often lack of coverage for particular features for all languages in the dataset, whereas the latter typically suffer from insufficient annotation and lack linguistic coverage. Both kinds of datasets have a particular kind of value, but do not go far enough to address the concerns of linguistic diversity (i.e. Ben´ıtez-Burraco 2025) and gradience (i.e. Levshina 2023) needed to expand our understanding of (cross-)linguistic processes.
@@
More recent attempts to create annotated databases that do address these issues by developing robust crosslinguistic corpora include the Universal Dependencies Treebanks project (UDT; Zeman et al. 2024), [^8] https://universaldependencies.org which contains data from individual languages that has been annotated for parts of speech and dependencies using a common framework. These kinds of datasets are promising in allowing for the automatic identification of properties that can categorize or classify languages in various types, which reflects an increasing interest in computational typology (Jäger 2025), whereby annotated corpora can allow for extraction of features that align with those found in typological databases. [^9] It should be noted here that so far the focus in computational typology has been less on automatically extracting features from corpora and more on validating theories of language/universals using expert classifications found in various descriptive typological databases, a consideration that I return to in the final discussion section below. Unfortunately, such datasets still suffer many of the same issues regarding coverage and degree of annotation as identified above.
@@
The amount of crosslinguistic coverage and degree of annotation of existing datasets are concerns when trying to investigate questions related to universal properties of human language. The reason for this is that such investigations require annotated data with information on word types or dependencies, data unavailable for most of the world’s languages. This means that crosslinguistic investigations using smaller datasets may suffer from lack of generalizability. For example, while Hahn & Xu (2022) investigated dependencies in 80 languages (the largest sample of languages to date for this kind of investigation), [^10] Recent cross-linguistic investigations using tagged corpora include Jing et al. (2021: 71 languages) and Hahn et al. (2020: 51 languages). the majority of languages in their dataset were from the Indo-European language family. This calls into question the generalizability of their findings, as such findings might be biased toward a particular group of languages (Blasi et al. 2022). [^11] Similar to how other scientific findings have been biased toward WEIRD demographics (Henrich et al. 2010).
@@
In its most recent form, the UDT has 160+ hand-annotated languages available for crosslinguistic investigation, but there are some concerns with the data. For one thing, the datasets are not parallel texts, which means there is a variety of genres representing each language, such that one could question how representative they are of a particular language as well as how easily they can be compared with corpora from other languages. Additionally, the number of sentences available for each language is variable, with many languages being represented by fewer than 200 sentences in the v2.14 release. Other datasets of tagged corpora have similar issues in terms of coverage. For example, the ODIN [^12] https://depts.washington.edu/uwcl/odin/ project (and its re-packaged form via LECS) [^13] https://huggingface.co/datasets/lecslab/glosslm-corpus contains interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences.
@@
These datasets and related publications show the value of annotated data for crosslinguistic investigation, but they also highlight some of the difficulties inherent in creating such data. For one thing, annotated data requires a large investment in time and manpower, whereby experts in languages must invest time to do actual annotation of the data. Additionally, annotators and researchers must decide which features should be annotated and to what degree. This is also partly dictated by the specifics of a language (which features are relatively more rare/interesting and/or well-described) as well as the amount, quality, and kind of descriptions available for a given language. Generally the data is sourced with a particular goal or research question in mind, which then conditions the manner and kind of annotation that is done.
@@
Recent advances in computational tools suggest the possibility of using such tools for automatic annotation, greatly reducing the time and effort needed to develop multilingual tagged datasets. As an example, there is some indication that automated taggers for under-resourced languages can benefit from high-resource language part of speech tags in parallel corpora (Agic et al., 2015; Imani et al., 2022b), particularly for major word classes like nouns and verbs, even in zero-shot situations on low-resource languages. But while the use of such tools may greatly reduce the annotation burden for researchers, the derived dataset must also be validated for crosslinguistic investigation. Specifically, it is incumbent on researchers to identify a set of criteria used to evaluate whether the automatic annotations are sufficiently aligned with human-curated annotations or classifications before such a dataset can be considered robust enough for further crosslinguistic investigation.
@@
The current paper reports on the development of a large corpus of parallel texts tagged for word class that partly addresses the issues identified above. With data sourced from the Parallel Bible Corpus (PBC; Mayer & Cysouw 2014: 1,597 languages) and additional data sourced from newly-available Bible translations (345 languages), computational word alignment tools allow for automatic tagging of word classes in 1,942 languages (roughly 25% of the world’s languages) that represent 155 families and 78 isolates.141414‘Isolates’ is defined here as languages found in the corpus that do not share family membership with another language in the corpus. The automated part of speech tagging method outlined in this paper was validated by comparing the results with output from existing automatic taggers for high-resource languages (SpaCy, Trankit) [^15] https://spacy.io/ ; http://nlp.uoregon.edu/trankit as well as with gold-standard tagged data in the UDT. The results indicate good alignment with both kinds of evaluation, providing a baseline from which to conduct crosslinguistic investigations. Further, features extracted from this dataset align with word order classifications in typological databases.
@@
Statistical properties of the derived corpus also comport with expectations regarding normality of language data. Importantly, this is partly because each language also has a large amount of data available for comparison. Individual corpora in the majority of these languages (over 1,850) contain more than 1,800 parallel sentences/verses, and the smallest corpus contains more than 700 sentences/verses. This dwarfs both the linguistic data and the language diversity available from other sources and the parallel nature of the data means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the taggedPBC) as well as explanations regarding its development and Python code for replicating the results described here, can be found at the linked repository. [^16] https://github.com/lingdoc/taggedPBC
@@
In the sections below I first outline the methodology used to develop the auto-taggers for each language (§2). I then highlight statistics of the dataset (§3) and describe tests used to assess tag validity (§3.1). I also show that this dataset allows for extraction of features associated with typological properties, identifying a novel measure derived from this dataset (the “N1 ratio”) that correlates with intransitive word order classifications in multiple typological databases (§3.2). I conclude the paper with discussion of ways the dataset could be improved to support more robust crosslinguistic investigations (§4.1) as well as potential avenues of research made possible by this dataset (§4.2).
@@ 2 Methodology
Transferring POS tags from high-resource languages has been shown to work relatively well for word classes with many members, such as Nouns and Verbs (Agic et al. 2015; Imani et al. 2022b, a). Specifically, Agic et al. (2015) show that using an IBM Model 2 to transfer tags between aligned words in pairs of languages achieves much better results than unsupervised and weakly-supervised approaches. Additionally, Imani et al. (2022b), using the PBC, were able to improve on accurate POS-tagging for low-resource languages by propagating labels from 36 high-resource languages to 17 low-resource target languages. They note that
@@
While Imani et al. (2022b)’s approach does result in higher overall accuracy than that of Agic et al. (2015) for specific language pairs, it is not clear that it performs better for particular word classes or for all low-resource languages more generally. Additionally, the computational resources required by Imani et al’s approach (using transformer architectures and word embeddings) are much more significant than those required by Agic et al.’s approach. Accordingly, the methodolgy used here follows that of Agic et al. (2015) in training IBM Model 2 to align words between language pairs, in this case each source language and English. It should be noted here that IBM Model 2 is a statistical word alignment model that is agnostic to the order of words, which makes it ideal for a situation like this one, where we are dealing with languages for which we don’t know the word order. The essential idea is that for each language pair we train a model to align words in a parallel corpus with two languages, then we “translate” the words in the source (low-resource) language to the target (high-resource) language, and transfer the POS-tag of the target language word to the source language word.
Another important thing to note is that while machine translation methods have expanded in recent years, there are various issues with some of the more recent (“AI”-related) approaches, particularly the use of large language models. Training or fine-tuning such models to translate between languages is not only computationally resource-intensive, but often also result in inaccuracies in output, making it questionable to transfer POS-tags. Further, developments in word alignment algorithms such as eflomal (Östling & Tiedemann 2016) and Embedding-Enhanced GIZA++ (Marchisio et al. 2022) require language-specific processing (embeddings in some cases) and can also be computationally resource-intensive. Additionally, alignment algorithms such as eflomal do not handle language pairs with word order differences particularly well. In contrast, the IBM Model 2 has relatively minor computational resource requirements and is simple to implement, while depending primarily on statistical patterns that handle differences in word order between language pairs.
@@ 2.1 Improving tag accuracy
Accordingly, to address the concerns present under the first principle, I reduce the number of verses in the corpus using the following criteria. I first lemmatize two English New Testament translations, then reduce the dataset to verses which share at least 4 lemmas in those translations. This ensures that verses are more likely to be semantically accessible. Translations chosen for this comparison are the New Living Translation (NLT) and the New International Version (NIV), after discussions with Bible translators who work with low-resource languages. Translators indicated that these two versions are often consulted for their readability and accessibility in terms of how they translate the original Greek text, particularly in backtranslation (a checking step before final translation verification). The smaller dataset is then further reduced to verses where a verb occurs that is also present in at least five other verses, which increases the likelihood that each individual non-English Bible translation has common concepts that can be aligned with words in the English text. This procedure results in 1,884 verses for training word alignment models. As noted above, this is a principled method for under-sampling the parallel data, which should help improve the quality of the trained statistical word alignment model.
@@
It should also be noted that the languages in the PBC use several different writing systems and scripts. Some, like Devanagari, are syllabic, while others, like Arabic, are largely phonetic, and yet others, like Chinese, are systems that use individual characters to represent words or syllables. In order to ensure that we can adequately compare languages in terms of their sound structure, and to support the training of subword tokenizers, I convert all disparate systems to romanized script using the uroman [^17] https://www.isi.edu/~ulf/uroman.html tool, with the additional step of performing word-level tokenization for translations using scripts where words are not clearly separated by whitespace (fugashi [^18] https://github.com/polm/fugashi for Japanese, khmernltk [^19] https://github.com/VietHoang1512/khmer-nltk for Khmer/Kuy, jieba [^20] https://github.com/fxsjy/jieba for Chinese, pyidaungsu [^21] https://github.com/kaunghtetsan275/pyidaungsu for Burmese/Karen/Mon, attacut [^22] https://github.com/PyThaiNLP/attacut for Thai, botok [^23] https://github.com/OpenPecha/Botok for Tibetan) before converting to romanized script. [^24] As noted by a reviewer, this is not an ideal mapping, since romanization does not necessarily reflect the phonemic values of a language in the same way that a standardized orthography does. Putting aside the question of phonemic or phonetic accuracy, however, I would argue that romanization does at least allow for comparison in a way that retaining original scripts does not.

### v3

@@ Abstract
Existing datasets available for crosslinguistic investigations have tended to focus on large amounts of data for a small group of languages or a small amount of data for a large number of languages. This means that claims based on these datasets are limited in what they reveal about universal properties of the human language faculty. While this has begun to change through the efforts of projects seeking to develop tagged corpora for a large number of languages, such efforts are still constrained by limits on resources. The current paper reports on a large tagged parallel dataset which has been developed to partially address this issue. The taggedPBC contains POS-tagged parallel text data from more than 1,940 languages, representing 155 language families and 78 isolates, dwarfing previously available resources. The accuracy of particular tags in this dataset is shown to correlate well with existing SOTA taggers for high-resource languages (SpaCy, Trankit). Additionally, a novel measure derived from this dataset, the N1 ratio, correlates with expert determinations of intransitive word order in three typological databases (WALS, Grambank, Autotyp) such that a Gaussian Naive Bayes classifier trained on this feature can accurately identify basic intransitive word order for languages not in those databases. While much work is still needed to expand and develop this dataset, the taggedPBC is an important step to enable corpus-based crosslinguistic investigations, and is made available for research and collaboration via GitHub.
keywords:
datasets, POS-tagging, low resource languages, computational typology
@@ 1 Introduction
There are over 7,000 languages in the world, and such languages are rather diverse in terms of their typological features. At the same time, despite the overall diversity there are many similarities between languages, suggesting the possibility of universal properties that guide language evolution. These diversities and similarities have been identified and categorized in various ways by a variety of databases, most notably WALS [Dryer:2013ab], Grambank [Skirgard:2023ab], Autotyp [Bickel:2023aa], Glottolog [Hammarstrom:2024aa], and Lexibank [List:2022aa]. [^1] See http://wals.info ; http://grambank.clld.org ; http://github.com/autotyp/autotyp-data ; http://www.glottolog.org ; http://lexibank.clld.org There are also collections of corpora that are expansive in their coverage of individual or groups of languages yet focused on particular genres or domains (CHILDES, Brown, etc; and see the extensive corpora available via the CLARIN project). [^2] See http://childes.talkbank.org ; http://www.clarin.eu/resource-families The former tend to provide information about languages, while the latter tend to provide actual language data in various forms. The issue with the former is often lack of coverage for particular features for all languages in the dataset, whereas the latter typically suffer from insufficient annotation and lack linguistic coverage. Both kinds of datasets have a particular kind of value, but do not go far enough to address the concerns of linguistic diversity (i.e. [Benitez-Burraco:2025aa]) and gradience (i.e. [Levshina:2023ab]) needed to expand our understanding of (cross-)linguistic processes.
@@
More recent attempts to create annotated databases that do address these issues by developing robust crosslinguistic corpora include the Universal Dependencies Treebanks project (UDT; [Zeman:2024ab]), [^3] See http://universaldependencies.org which contains data from individual languages that has been annotated for parts of speech and dependencies using a common framework. These kinds of datasets are promising in allowing for the automatic identification of properties that can categorize or classify languages in various types, which reflects an increasing interest in computational typology [Jager:2025aa], whereby annotated corpora can allow for extraction of features that align with those found in typological databases. [^4] It should be noted here that so far the focus in computational typology has been less on automatically extracting features from corpora and more on validating theories of language/universals using expert classifications found in various descriptive typological databases, a consideration that I return to in the final discussion section below. Unfortunately, such datasets still suffer many of the same issues regarding coverage and degree of annotation as identified above.
@@
The amount of crosslinguistic coverage and degree of annotation of existing datasets are concerns when trying to investigate questions related to universal properties of human language. The reason for this is that such investigations require annotated data with information on word types or dependencies, data unavailable for most of the world’s languages. This means that crosslinguistic investigations using smaller datasets may suffer from lack of generalizability. For example, while [Hahn:2022aa] investigated dependencies in 80 languages (the largest sample of languages to date for this kind of investigation), [^5] Recent cross-linguistic investigations using tagged corpora include [Jing:2021aa] (71 languages) and [Hahn:2020aa] (51 languages). the majority of languages in their dataset were from the Indo-European language family. This calls into question the generalizability of their findings, as such findings might be biased toward a particular group of languages [Blasi:2022aa]. [^6] Similar to how other scientific findings have been biased toward WEIRD demographics [Henrich:2010aa].
@@
In its most recent form, the UDT has 160+ hand-annotated languages available for crosslinguistic investigation, but there are some concerns with the data. For one thing, the datasets are not parallel texts, which means there is a variety of genres representing each language, such that one could question how representative they are of a particular language as well as how easily they can be compared with corpora from other languages. Additionally, the number of sentences available for each language is variable, with many languages being represented by fewer than 200 sentences in the v2.14 release. Other datasets of tagged corpora have similar issues in terms of coverage. For example, the ODIN project (and its re-packaged form via LECS) [^7] See http://depts.washington.edu/uwcl/odin ; http://huggingface.co/datasets/lecslab/glosslm-corpus contains interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences.
@@
These datasets and related publications show the value of annotated data for crosslinguistic investigation, but they also highlight some of the difficulties inherent in creating such data. For one thing, annotated data requires a large investment in time and manpower, whereby experts in specific languages must invest time to do actual annotation of the data. Additionally, annotators and researchers must decide which features should be annotated and to what degree. This is also partly dictated by the specifics of a language (which features are relatively more rare/interesting and/or well-described) as well as the amount, quality, and kind of descriptions available for a given language. Generally the data is sourced with a particular goal or research question in mind, which then conditions the manner and kind of annotation that is done.
@@
Recent advances in computational tools suggest the possibility of using such tools for automatic annotation, greatly reducing the time and effort needed to develop multilingual tagged datasets. As an example, there is some indication that automated taggers for under-resourced languages can benefit from high-resource language part of speech tags in parallel corpora [Agic:2015aa, Imani:2022aa], particularly for major word classes like nouns and verbs, even in zero-shot situations on low-resource languages. But while the use of such tools may greatly reduce the annotation burden for researchers, the derived dataset must also be validated for crosslinguistic investigation. Specifically, it is incumbent on researchers to identify a set of criteria used to evaluate whether the automatic annotations are sufficiently aligned with human-curated annotations or classifications before such a dataset can be considered robust enough for further crosslinguistic investigation.
@@
The current paper reports on the development of a large corpus of parallel texts tagged for word class that partly addresses the issues identified above. With data sourced from the Parallel Bible Corpus (PBC; [Mayer:2014aa]: 1,597 languages) and additional data sourced from newly-available Bible translations (345 languages), computational word alignment tools allow for automatic tagging of word classes in 1,942 languages (roughly 25% of the world’s languages) that represent 155 families and 78 isolates.888‘Isolates’ is defined here as languages found in the corpus that do not share family membership with another language in the corpus. The automated part of speech tagging method outlined in this paper was validated by comparing the results with output from existing automatic taggers for high-resource languages (SpaCy, Trankit). [^9] See http://spacy.io/ ; http://nlp.uoregon.edu/trankit The results indicate good alignment, providing a baseline from which to conduct crosslinguistic investigations. Further, features extracted from this dataset align with word order classifications in typological databases.
@@
Statistical properties of the derived corpus also comport with expectations regarding normality of language data. Importantly, this is partly because each language has a large amount of data available for comparison. Individual corpora in the majority of these languages (over 1,850) contain more than 1,800 parallel sentences/verses, and the smallest corpus contains more than 700 sentences/verses. This dwarfs both the linguistic data and the language diversity available from other sources. Further, the parallel nature of the data means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the taggedPBC) as well as explanations regarding its development and Python code for replicating the results described here, can be found at the linked repository. [^10] See http://github.com/lingdoc/taggedPBC
@@
In the sections below I first outline the methodology used to develop the auto-taggers for each language (§2). I then highlight statistics of the dataset (§3) and describe the test used to assess tag validity (§3.1). I also show that this dataset allows for extraction of features associated with typological properties, identifying a novel measure derived from this dataset (the “N1 ratio”) that correlates with intransitive word order classifications in multiple typological databases (§3.2). I conclude the paper with discussion of ways the dataset could be improved to support more robust crosslinguistic investigations (§4.1) as well as potential avenues of research made possible by this dataset (§4.2).
@@ 2 Methodology
Transferring POS tags from high-resource languages has been shown to work relatively well for word classes with many members, such as Nouns and Verbs [Agic:2015aa, Imani:2022aa, Imani:2022ab]. Specifically, [Agic:2015aa] show that using an IBM Model 2 to transfer tags between aligned words in pairs of languages achieves much better results than unsupervised and weakly-supervised approaches. Additionally, [Imani:2022aa], using the PBC, were able to improve on accurate POS-tagging for low-resource languages by propagating labels from 36 high-resource languages to 17 low-resource target languages. They note that
@@
While [Imani:2022aa]’s approach does result in higher overall accuracy than that of [Agic:2015aa] for specific language pairs, it is not clear that it performs better for particular word classes or for all low-resource languages more generally. Additionally, the computational resources required by Imani et al’s approach (using transformer architectures and word embeddings) are much more significant than those required by [Agic:2015aa]’s approach. Accordingly, the methodolgy used here follows that of [Agic:2015aa] in training IBM Model 2 to align words between language pairs, in this case each source language and English. It should be noted here that IBM Model 2 is a statistical word alignment model that is agnostic to the order of words, which makes it ideal for a situation like this one, where we are dealing with languages for which we don’t know the word order. The essential idea is that for each language pair we train a model to align words in a parallel corpus with two languages, then we “translate” the words in the source (low-resource) language to the target (high-resource) language, and transfer the POS-tag of the target language word to the source language word.
Another important thing to note is that while machine translation methods have expanded in recent years, there are various issues with some of the more recent (“AI”-related) approaches, particularly the use of large language models. Training or fine-tuning such models to translate between languages is not only computationally resource-intensive, but often also result in inaccuracies in output, making it questionable to transfer POS-tags. Further, developments in word alignment algorithms such as eflomal [Ostling:2016aa] and Embedding-Enhanced GIZA++ [Marchisio:2022aa] require language-specific processing (embeddings in some cases) and can also be computationally resource-intensive. Additionally, alignment algorithms such as eflomal do not handle language pairs with word order differences particularly well. In contrast, the IBM Model 2 has relatively minor computational resource requirements and is simple to implement, while depending primarily on statistical patterns that handle differences in word order between language pairs.
@@ 2.1 Improving tag accuracy
Accordingly, to address the concerns present under the first principle, I reduce the number of verses in the corpus using the following criteria. I first lemmatize two English New Testament translations, then reduce the dataset to verses which share at least 4 lemmas in those translations. This ensures that verses are more likely to be semantically accessible. Translations chosen for this procedure are the New Living Translation (NLT) and the New International Version (NIV), after discussions with Bible translators who work with low-resource languages. Translators indicated that these two versions are often consulted for their readability and accessibility in terms of how they translate the original Greek text, particularly in backtranslation (a checking step before final translation verification). The smaller dataset is then further reduced to verses where a verb occurs that is also present in at least five other verses, which increases the likelihood that each individual non-English Bible translation has common concepts that can be aligned with words in the English text. This procedure results in 1,884 verses for training word alignment models. As noted above, this is a systematic method for under-sampling the parallel data, which should help improve the quality of the trained statistical word alignment model.
@@
It should also be noted that the languages in the PBC use several different writing systems and scripts. Some, like Devanagari, are syllabic, while others, like Arabic, are largely phonetic, and yet others, like Chinese, are systems that use individual characters to represent words or syllables. In order to ensure that we can adequately compare languages in terms of their sound structure, and to support the training of subword tokenizers, I convert all disparate systems to romanized script using the uroman [^11] See http://github.com/isi-nlp/uroman tool, with the additional step of performing word-level tokenization for translations using scripts where words are not clearly separated by whitespace (fugashi [^12] See http://github.com/polm/fugashi for Japanese, khmernltk [^13] See http://github.com/VietHoang1512/khmer-nltk for Khmer/Kuy, jieba [^14] See http://github.com/fxsjy/jieba for Chinese, pyidaungsu [^15] See http://github.com/kaunghtetsan275/pyidaungsu for Burmese/Karen/Mon, attacut [^16] See http://github.com/PyThaiNLP/attacut for Thai, botok [^17] See http://github.com/OpenPecha/Botok for Tibetan) before converting to romanized script. [^18] As noted by a reviewer, this is not an ideal mapping, since romanization does not necessarily reflect the phonemic values of a language in the same way that a standardized orthography does. Putting aside the question of phonemic or phonetic accuracy, however, I would argue that romanization does at least allow for comparison in a way that retaining original scripts does not.